A bus travels 60km due north in 1 hour.It then travels 75km due east in 2 hours. What is the average velocity of the bus


1
Expert's answer
2021-03-16T11:36:11-0400

By the definition of the average velocity, we have:


"v_{avg}=\dfrac{d_{tot}}{t_{tot}}."

Let's find the total time:


"t_{tot}=t_1+t_2=1\ h+2\ h=3\ h."


Let's write the resultant vector of the average velocity:


"v_{avg}=(\dfrac{75}{3})i+(\dfrac{60}{3})j."


We can find the magnitude of the total displacement from the Pythagorean theorem:


"d_{tot}=\sqrt{d_x^2+d_y^2}=\sqrt{(75\ km)^2+(60\ km)^2}=96\ km."

Therefore, the magnitude of the average velocity can be found as follows:


"v_{avg}=\dfrac{96\ km}{3\ h}=32\ \dfrac{km}{h}."

We can find the direction of the average velocity as follows:


"\theta=tan^{-1}(\dfrac{v_y}{v_x}),""\theta=tan^{-1}(\dfrac{60\ km}{75\ km})=38.6^{\circ}\ N\ of\ E."